OpenStudy (anonymous):

(1over4(x-5)+(5over8)=5x-(5over8)

OpenStudy (anonymous):

please improve the description of the equation

OpenStudy (agreene):

\[\frac{1}{4(x-5)}+\frac58=5x-\frac58\] is this your question?

OpenStudy (anonymous):

1/4(x-5)+5/8=5x-5/8

OpenStudy (anonymous):

thanks, infinitely more helpful

OpenStudy (agreene):

\[\frac14 (x-5)\] \[\frac{1}{4(x-5)}\] which of these is the first term? both times you wrote it its hard to tell.

OpenStudy (anonymous):

the top one .25(x-5)

OpenStudy (agreene):

x=0 (x-5)/4 +5/8 = 5x-5/8 x/4-5/8=5x-5/8 add 5/8-5x to both sides -19x/4= 0 x=0

OpenStudy (anonymous):

thank you